Sunshine Tours is a travel agency specializing in flights between Toronto and Jamaica.




It books passenger on Canadian air and charges passengers $1,000 per round-trip ticket.




Sunshine’s fixed costs are $22,000 per month. Its variable costs per ticket, including a delivery




fee is $18 (assume each ticket is purchased in a separate package. Thus the delivery fee applies




to each ticket)




Maryland International College Page 3




Required: -




1) What number of tickets Sunshine must sell each month to




a) Break even




b) Make a target operating income of $10,000




2) Assume another company, TNT Express, offers to charge Sunshine only $12 per ticket.




How would accepting this offer affect your answer to (a) and (b) in requirement 1?

1
Expert's answer
2022-01-18T09:34:01-0500

1)a)

"Q=\frac{FC}{P-VC}=\frac{22000}{1000-36}=22.82"



b)"Q=\frac{10000}{10000-36}=10.37"



2)"Q=\frac{FC}{P-VC}=\frac{22000}{1000-24}=22.54"



A similar level of contribution margin can be achieved at a lower variable cost price